Let α, β be roots of the equation x^2− 3x − 1 = 0. For each nonnegative integer n,

let y_n = α^n + β^n

. Show that gcd(y_n, y_(n+1)) = 1 for each nonnegative integer n.


1
Expert's answer
2021-11-03T17:13:22-0400

Let's consider the equation "x^2-3x-1=0." The discriminant equals to "D=9+4=13". So the equation has irrational roots: "\alpha=\frac{3+√3}{2}, \beta=\frac{3-√3}{2}" . From the Viet theorem "\alpha+\beta=3, \alpha\beta=-1." If "n=0" we have "y_0=\alpha^0+\beta^0=2" . If "n=1" we have "y_1=\alpha+\beta=3." So "gcd(y_0;y_1)=gcd(2;3)=1." In case "n=2" we have "y_2=(\alpha+\beta)^2-2\alpha\beta=9+2=11, gcd(y_1;y_2)=1" Let's prove that "y_n" is natural and "gcd(y_n, y_{n+1})=1" with the help of the method of mathematical induction. The base of induction is true. Let the proposition is true for all "n=1,...,k\in N" , and "y_1,y_k \in N. gcd(y_{n-1},y_n)=1." If "n=k+1" we have "y_{n+1}=(\alpha+\beta)y_n+\alpha\beta y_{n-1}" . Si we have "y_{n+1} \in N." And "gcd(y_k, y_{k+1})=gcd(y_k,y_{k-1})=1."